# Beyond MMD: Evaluating Graph Generative Models with Geometric Deep Learning

### v1

@@ I Introduction
Complex systems, characterized by many non-trivially interacting components, are pervasive in nature, society, and technology and often exhibit emergent phenomena, self-organization, and complex dynamics. Researchers often represent them as networks – graphs in the mathematical sense – where nodes (vertices) represent entities and links (edges) represent relationships between them. Due to their characteristics, these networks are often referred to as complex networks, and a new field of research, called Network Science, has emerged to study them. To understand and analyze complex networks and the underlying systems, researchers rely on mathematical, physical and computational models, statistics, and machine learning techniques. For instance, the degree distribution, clustering coefficient, and average path length are some of the most common topological properties used to characterize complex networks, while the presence of motifs, community structure, hierarchical organization, and assortativity (i.e., the tendency of nodes to connect to similar nodes, homophily in jargon) are also important features that can provide insights into the underlying processes and mechanisms that govern the network’s behavior [newman2018networks]. Moreover, robustness, resilience, and vulnerability are also important properties to consider when analyzing complex networks, as they can provide insights into the network’s ability to withstand perturbations or attacks [grassia2024robustness, grassia2021machine]. An open key question in Network Science is how to model real-world networks that capture all these properties, as they are often hard to quantify and reproduce. In fact, the ability to generate realistic networks is crucial in many fields including bioinformatics, social sciences, (network) medicine, and others, since it enables the evaluation of graph algorithms, the study of network dynamics, and the generation of null models.
@@
Graph generation is the process of creating synthetic graphs that have desired properties, ultimately mimicking the characteristics of real-world networks. While a plethora of non-trainable graph generation models have been proposed, such as the Erdős-Rényi model [erdos59a], the Barabási-Albert model [Baraba_si_1999], and the Stochastic Block Model [HOLLAND1983109], they reflect only a limited set of graph properties — such as degree distribution or clustering coefficient, or a trivial combination of them — that are insufficient to capture the complexity of real-world graphs and their underlying systems. This often results in unrealistic structures with limited applicability. For example, the Erdős-Rényi model generates graphs with a Poisson degree distribution, while the Barabási-Albert model produces scale-free networks with a power-law degree distribution thanks to its preferential attachment mechanism. This process favors the growth of hubs and leads to the emergence of a few highly connected nodes due to the “rich get richer” phenomenon. However, these models fail to capture other important properties of real-world networks, such as community structure, clustering, and the presence of motifs. The Stochastic Block Model can generate graphs with a predefined community structure, but it often relies on strong assumptions about the underlying graph structure and the number of communities, which may not be fixed in real-world networks. This limitation is partially overcome, for instance, by the LFR (Lancichinetti-Fortunato-Radicchi) benchmark [Lancichinetti_2008], a widely used synthetic graph generator that creates graphs with a power-law degree and community distributions, but it is still limited in its ability to capture other important properties, such as clustering and motifs. Other generative models, such as the nPSO (non-uniform Popularity-Similarity Optimization) model [Muscoloni_2018], sample from a non-uniform distribution of angular coordinates in the hyperbolic disk — typically a Gaussian mixture distribution — to create agglomerates of nodes that are more densely connected within their communities than with the rest of the network, and also allow controlling the clustering coefficient. However, all these algorithms inherently encode human assumptions about the underlying graph structure, which might work well for specific applications but fail to generalize across different domains, making them inadequate for applications that demand a more realistic representation of complex systems. On the other hand, in the last decade, the use of machine learning (and deep learning) has revolutionized many domains, including computer vision, natural language processing, and speech recognition, where deep learning techniques have proven effective in capturing the complex and diverse characteristics of real-world data. For instance, deep learning has been successfully used in addressing open questions in complex systems like the prediction of protein structures [article], the generation of realistic images [goodfellow2014generativeadversarialnetworks], and approaching combinatorial and NP–hard problems [10.1145/3551642, 10.1371/journal.pone.0296185]. In the context of graph generation, Graph Generative Models (GGMs) have emerged as a promising data-driven solution to this problem, leveraging deep learning techniques to learn the underlying distribution of real-world graphs and generate new samples that closely resemble them. Examples of GGMs include Variational Graph Autoencoders (VGAE) [kipf2016variationalgraphautoencoders], Graph Recurrent Attention Networks (GRAN), and more recently diffusion-based models like “Efficient and Degree-guided graph Generative model” (EDGE). However, the main limitation of these models often lies in the evaluation process, which typically relies on Maximum Mean Discrepancy (MMD) as a metric, which is not suitable for properly assessing the quality of generated graphs.
@@ II-A Graph Generation
The ER model [erdos59a] is the first ever proposed algorithm for the formation of random graphs. In its standard form, denoted as $G(n,p)$, it generates a graph with $n$ nodes where each possible edge between a pair of nodes is independently included with probability $p$. This model leads to graphs that are typically homogeneous, with node degrees concentrated around their mean, and no inherent geometric or community structure. ER graphs do not exhibit community structure and have a homogeneous degree distribution (meaning most nodes have similar degrees), which results in robustness to random failures and to targeted attacks. Yet, the Erdős-Rényi model is well known for its simplicity and efficiency in generating random graphs, has been widely studied, and serves as a foundational model in graph theory. For such reasons, it is often used, for instance as a null model for comparison with real-world networks, as it provides a baseline for understanding the significance of observed network properties.
@@
The SBM [HOLLAND1983109] is a generative model for random graphs that captures community structure by dividing nodes into blocks/groups and assigning edge probabilities based on block memberships. That is, the probability of an edge between two nodes depends on the blocks to which they belong. Formally, given a set of nodes partitioned into $K$ blocks, edges between node pairs are drawn independently according to a block-dependent probability matrix $P\in[0,1]^{K\times K}$, where $P_{ij}$ defines the probability of an edge between a node in block $i$ and a node in block $j$. It can be viewed as a generalization of the Erdős-Rényi model, where the probability of edge formation is not uniform across all nodes. The SBM model is particularly useful for modeling networks with community structure, where nodes within the same block are more likely to be connected than nodes in different blocks, but it suffers from limitations such as the need to specify the number of blocks and their sizes in advance, which may not be known in practice.
@@
The Barabási–Albert model [Baraba_si_1999] constructs a growing graph by simulating the process of preferential attachment, where new nodes are added to the graph and preferentially attach to existing nodes with a probability proportional to their degree, giving more connected nodes a higher chance of receiving new links in a “rich-get-richer” fashion. In particular, the model starts with a small number of nodes connected with arbitrary topology (often a fully connected graph), and iteratively adds new nodes, where each new node connects to a fixed number $m$ of existing nodes with probability proportional to their degrees. Formally, the probability $p_{i}$ that the new node is connected to node $i$ is:
@@
The LFR benchmark [Lancichinetti_2008] is a widely used synthetic graph generator that creates graphs with a power-law degree and community distributions. It often serves as a benchmark for testing and comparing community detection algorithms, as it allows researchers to generate networks with known community structures and evaluate the performance of their algorithms in detecting these communities. The LFR model includes a mixing parameter $\mu$ that controls the fraction of edges that are between communities, reflecting the amount of “noise” in the network: $\mu=0$ indicates no inter-community edges and $\mu=1$ indicates all edges are inter-community. The graph is generated by sampling a degree sequence from a power-law distribution with a specified exponent, assigning nodes to communities such that each community size follows another power-law distribution, while also considering the limitations imposed by the mixing parameter $\mu$, and finally adding links between nodes based on their community memberships and the mixing parameter.
@@
The nPSO [Muscoloni_2018] model is designed to generate complex networks with realistic community structures by addressing the limitations of previous models like PSO [Papadopoulos_2012] and GPA [Zuev_2015]. In the original Popularity-Similarity-Optimization (PSO) model, a graph is generated by simulating its growth process in a hyperbolic space, where nodes are placed in a polar coordinate system and connected based on their distance from each other. Specifically, two factors are considered: Popularity, which is represented by the radial coordinate of the nodes, and Similarity, represented by the angular coordinate. During growth, new nodes are added to the graph and connected to a fixed number of existing nodes based on their distance in the hyperbolic space, which is iteratively increased to simulate the fading of popularity. The Geometrical-Preferential-Attachment (GPA) model builds on the PSO model by introducing a third factor, called Attractiveness, a form of preferential attachment based on hyperbolic distance between nodes, that determines the probability of connecting to existing nodes and influences the growth process and the resulting graph structure, which exhibits scale-free degree distribution and a non-trivial community structure. nPSO further refines the model by sampling from a non-uniform distribution of nodes over the hyperbolic disk, which allows for the creation of agglomerates of nodes that are more densely connected within their communities than with the rest of the network. For instance, if a Gaussian mixture distribution is used, the resulting graph will exhibit communities that emerge in correspondence with the different Gaussians, allowing the control of the number of communities and their sizes.
@@ II-B Graph Classification and Graph Similarity
Graph classification is the task of assigning a label to a graph based on its structure and properties. Formally, given a set of graphs $G=\{G_{1},\ldots,G_{N}\}$ and their labels $y=\{y_{1},\ldots,y_{N}\}$, the goal is to predict the label $y_{i}$ of each graph $G_{i}$. Many approaches have been proposed, including traditional methods that are based on hand-crafted feature extraction and more recent deep learning-based ones. For instance, traditional methods often rely on graph kernels, which first extract a set of hand-crafted features (e.g., node degree, clustering coefficient, node coloring, etc.) and then use a kernel function to measure the similarity between graphs based on these features. On the other hand, deep learning-based methods leverage neural networks to learn representations of graphs, allowing for more flexible and powerful classification models, which may not involve hand-crafted features or some possibly biased assumptions about the graph structure, as it learns the features directly from the data. In this context, Graph Representation Learning (GRL), also known as Geometric Deep Learning [DBLP:journals/corr/BronsteinBLSV16], has emerged as a powerful framework for learning representations of graph-structured data. The basic idea is that — as common in most deep learning approaches — the model learns a mapping function from the input space (i.e., the graph) to a high-dimensional embedding space (i.e., a latent space where a graph $G_{i}$ is represented as a vector $h_{G_{i}}$) using a neural network. Such representation should preserve some useful properties of the original input data, and should be useful for the downstream task like, but not limited to, graph classification.
@@
The most popular approach to learn representations of graphs are Graph Neural Networks (GNNs) [hamilton2020graph, bronstein2021geometric, 4700287], based on a message-passing mechanism that allows each node to exchange information with its neighbors and to build its own representation based on a trainable function of the information received from its neighbors. That is, GNNs iteratively update the representation of each node by aggregating incoming messages from its neighbors, which is, in turn, sent to its own neighbors in the next iteration to update their representations. A plethora of GNNs have been proposed in the literature, each with its own architecture and design choices. For instance, Graph Convolutional Networks (GCNs) [kipf2017semisupervisedclassificationgraphconvolutional] use a convolutional operator (i.e., a shared filter) to aggregate information from neighboring nodes, GraphSAGE [hamilton2018inductiverepresentationlearninglarge] extends this idea by using a sampling strategy and sampling a fixed-size neighborhood to allow the model to scale to large graphs, while Graph Attention Networks (GATs) [veličković2018graphattentionnetworks, grassia2021wsgat] use an attention mechanism to weigh the importance of neighboring nodes when aggregating their features, allowing the model to focus on the most relevant ones for the task, and Graph Isomorphism Networks (GIN) [DBLP:journals/corr/abs-1810-00826] are inspired by the Weisfeiler-Lehman graph isomorphism test and use a neural network (e.g., a multi-layer perceptron) to learn a representation of the graph based on its structure.
@@
From node representations, graph-level representations (i.e., the graph embedding) can be obtained by pooling the representation of its most elementary components (i.e., nodes) using a permutation-invariant pooling function like global pooling functions (e.g., mean, max, sum) or by leveraging more sophisticated techniques such as attention mechanisms or hierarchical aggregation [ying2019hierarchicalgraphrepresentationlearning, ranjan2020asapadaptivestructureaware]. Specifically, a GDL model is usually composed of a series of GNN layers, interspersed with non-linear activation functions, followed by the pooling layer. That is, $h_{G}=\operatorname{pool}(h_{v_{1}},h_{v_{2}},\ldots,h_{v_{n}})$, where $h_{G}$ is the graph-level representation, $h_{v_{i}}$ is the representation of node $v_{i}$, and $\operatorname{pool}$ is a permutation-invariant pooling function.
@@
Once the graph-level representation is obtained, it can be used for various tasks (including, of course, graph classification) using classic ML techniques, such as fully connected layers, support vector machines, or other classifiers. In this sense, similarity between two graphs can be defined as the distance between their respective representations in the embedding space, which is often learned by the GDL model for a specific task, such as graph classification. More formally, given two graphs $G_{i}$ and $G_{j}$, they are similar if their representations $h_{G_{i}}$ and $h_{G_{j}}$ are close in the embedding space. In this context, graph similarity computation [jin2022cgmncontrastivegraphmatching, DBLP:journals/corr/abs-1912-11615] has seen significant advancements in recent years and has proven to be highly beneficial across various domains, but has not yet been tested in the field of graph generation extensively. A method for improving graph generation evaluation using a classifier was proposed in the literature by Liu et al. [Liu2019AutoregressiveGG], although it was demonstrated on a limited dataset and with restricted use of Graph Generative Models. Specifically, they trained both the Graph Generative Models and a Graph Isomorphism Network (GIN) classifier on small graph datasets such as Community-small, Grid-small, and Ego-small.While this work represents a step forward, it has several limitations, including the use of small datasets and the limited evaluation of the generated graphs. Here, we address these limitations by proposing a more comprehensive evaluation methodology that leverages larger datasets that include both real-world and synthetic graphs, by employing a more sophisticated model architecture and training methodology, and by providing a more thorough evaluation of the generated graphs.

### v2

@@ I Introduction
Complex systems, characterized by many non-trivially interacting components, are pervasive in nature, society, and technology and often exhibit emergent phenomena, self-organization, and complex dynamics. Researchers often represent them as networks – graphs in the mathematical sense – where nodes (vertices) represent entities and links (edges) represent relationships between them. Due to their characteristics, these networks are often referred to as complex networks, and a new field of research, called Network Science, has emerged to study them. To understand and analyze complex networks and the underlying systems, researchers rely on mathematical, physical and computational models, statistics, and machine learning techniques. For instance, the degree distribution, clustering coefficient, and average path length are some of the most common topological properties used to characterize complex networks, while the presence of motifs, community structure, hierarchical organization, and assortativity (i.e., the tendency of nodes to connect to similar nodes, homophily in jargon) are also important features that can provide insights into the underlying processes and mechanisms that govern the network’s behavior [34]. Moreover, robustness, resilience, and vulnerability are also important properties to consider when analyzing complex networks, as they can provide insights into the network’s ability to withstand perturbations or attacks [1, 13]. An open key question in Network Science is how to model real-world networks that capture all these properties, as they are often hard to quantify and reproduce. In fact, the ability to generate realistic networks is crucial in many fields including bioinformatics, social sciences, (network) medicine, and others, since it enables the evaluation of graph algorithms, the study of network dynamics, and the generation of null models.
@@
Graph generation is the process of creating synthetic graphs that have desired properties, ultimately mimicking the characteristics of real-world networks. While a plethora of non-trainable graph generation models have been proposed, such as the Erdős-Rényi model [11], the Barabási-Albert model [2], and the Stochastic Block Model [23], they reflect only a limited set of graph properties — such as degree distribution or clustering coefficient, or a trivial combination of them — that are insufficient to capture the complexity of real-world graphs and their underlying systems. This often results in unrealistic structures with limited applicability. For example, the Erdős-Rényi model generates graphs with a Poisson degree distribution, while the Barabási-Albert model produces scale-free networks with a power-law degree distribution thanks to its preferential attachment mechanism. This process favors the growth of hubs and leads to the emergence of a few highly connected nodes due to the “rich get richer” phenomenon. However, these models fail to capture other important properties of real-world networks, such as community structure, clustering, and the presence of motifs. The Stochastic Block Model can generate graphs with a predefined community structure, but it often relies on strong assumptions about the underlying graph structure and the number of communities, which may not be fixed in real-world networks. This limitation is partially overcome, for instance, by the LFR (Lancichinetti-Fortunato-Radicchi) benchmark [28], a widely used synthetic graph generator that creates graphs with a power-law degree and community distributions, but it is still limited in its ability to capture other important properties, such as clustering and motifs. Other generative models, such as the nPSO (non-uniform Popularity-Similarity Optimization) model [33], sample from a non-uniform distribution of angular coordinates in the hyperbolic disk — typically a Gaussian mixture distribution — to create agglomerates of nodes that are more densely connected within their communities than with the rest of the network, and also allow controlling the clustering coefficient. However, all these algorithms inherently encode human assumptions about the underlying graph structure, which might work well for specific applications but fail to generalize across different domains, making them inadequate for applications that demand a more realistic representation of complex systems. On the other hand, in the last decade, the use of machine learning (and deep learning) has revolutionized many domains, including computer vision, natural language processing, and speech recognition, where deep learning techniques have proven effective in capturing the complex and diverse characteristics of real-world data. For instance, deep learning has been successfully used in addressing open questions in complex systems like the prediction of protein structures [25], the generation of realistic images [12], and approaching combinatorial and NP–hard problems [6, 7]. In the context of graph generation, Graph Generative Models (GGMs) have emerged as a promising data-driven solution to this problem, leveraging deep learning techniques to learn the underlying distribution of real-world graphs and generate new samples that closely resemble them. Examples of GGMs include Variational Graph Autoencoders (VGAE) [26], Graph Recurrent Attention Networks (GRAN), and more recently diffusion-based models like “Efficient and Degree-guided graph Generative model” (EDGE). However, the main limitation of these models often lies in the evaluation process, which typically relies on Maximum Mean Discrepancy (MMD) as a metric, which is not suitable for properly assessing the quality of generated graphs.
@@ II-A Graph Generation
The ER model [11] is the first ever proposed algorithm for the formation of random graphs. In its standard form, denoted as $G(n,p)$, it generates a graph with $n$ nodes where each possible edge between a pair of nodes is independently included with probability $p$. This model leads to graphs that are typically homogeneous, with node degrees concentrated around their mean, and no inherent geometric or community structure. ER graphs do not exhibit community structure and have a homogeneous degree distribution (meaning most nodes have similar degrees), which results in robustness to random failures and to targeted attacks. Yet, the Erdős-Rényi model is well known for its simplicity and efficiency in generating random graphs, has been widely studied, and serves as a foundational model in graph theory. For such reasons, it is often used, for instance as a null model for comparison with real-world networks, as it provides a baseline for understanding the significance of observed network properties.
@@
The SBM [23] is a generative model for random graphs that captures community structure by dividing nodes into blocks/groups and assigning edge probabilities based on block memberships. That is, the probability of an edge between two nodes depends on the blocks to which they belong. Formally, given a set of nodes partitioned into $K$ blocks, edges between node pairs are drawn independently according to a block-dependent probability matrix $P\in[0,1]^{K\times K}$, where $P_{ij}$ defines the probability of an edge between a node in block $i$ and a node in block $j$. It can be viewed as a generalization of the Erdős-Rényi model, where the probability of edge formation is not uniform across all nodes. The SBM model is particularly useful for modeling networks with community structure, where nodes within the same block are more likely to be connected than nodes in different blocks, but it suffers from limitations such as the need to specify the number of blocks and their sizes in advance, which may not be known in practice.
@@
The Barabási–Albert model [2] constructs a growing graph by simulating the process of preferential attachment, where new nodes are added to the graph and preferentially attach to existing nodes with a probability proportional to their degree, giving more connected nodes a higher chance of receiving new links in a “rich-get-richer” fashion. In particular, the model starts with a small number of nodes connected with arbitrary topology (often a fully connected graph), and iteratively adds new nodes, where each new node connects to a fixed number $m$ of existing nodes with probability proportional to their degrees. Formally, the probability $P_{i}$ that the new node is connected to node $i$ is:
@@
The LFR benchmark [28] is a widely used synthetic graph generator that creates graphs with a power-law degree and community distributions. It often serves as a benchmark for testing and comparing community detection algorithms, as it allows researchers to generate networks with known community structures and evaluate the performance of their algorithms in detecting these communities. The LFR model includes a mixing parameter $\mu$ that controls the fraction of edges that are between communities, reflecting the amount of “noise” in the network: $\mu=0$ indicates no inter-community edges and $\mu=1$ indicates all edges are inter-community. The graph is generated by sampling a degree sequence from a power-law distribution with a specified exponent, assigning nodes to communities such that each community size follows another power-law distribution, while also considering the limitations imposed by the mixing parameter $\mu$, and finally adding links between nodes based on their community memberships and the mixing parameter.
@@
The nPSO [33] model is designed to generate complex networks with realistic community structures by addressing the limitations of previous models like PSO [36] and GPA [44]. In the original Popularity-Similarity-Optimization (PSO) model, a graph is generated by simulating its growth process in a hyperbolic space, where nodes are placed in a polar coordinate system and connected based on their distance from each other. Specifically, two factors are considered: Popularity, which is represented by the radial coordinate of the nodes, and Similarity, represented by the angular coordinate. During growth, new nodes are added to the graph and connected to a fixed number of existing nodes based on their distance in the hyperbolic space, which is iteratively increased to simulate the fading of popularity. The Geometrical-Preferential-Attachment (GPA) model builds on the PSO model by introducing a third factor, called Attractiveness, a form of preferential attachment based on hyperbolic distance between nodes, that determines the probability of connecting to existing nodes and influences the growth process and the resulting graph structure, which exhibits scale-free degree distribution and a non-trivial community structure. nPSO further refines the model by sampling from a non-uniform distribution of nodes over the hyperbolic disk, which allows for the creation of agglomerates of nodes that are more densely connected within their communities than with the rest of the network. For instance, if a Gaussian mixture distribution is used, the resulting graph will exhibit communities that emerge in correspondence with the different Gaussians, allowing the control of the number of communities and their sizes.
@@ II-B Graph Classification and Graph Similarity
Graph classification is the task of assigning a label to a graph based on its structure and properties. Formally, given a set of graphs $G=\{G_{1},\ldots,G_{N}\}$ and their labels $y=\{y_{1},\ldots,y_{N}\}$, the goal is to predict the label $y_{i}$ of each graph $G_{i}$. Many approaches have been proposed, including traditional methods that are based on hand-crafted feature extraction and more recent deep learning-based ones. For instance, traditional methods often rely on graph kernels, which first extract a set of hand-crafted features (e.g., node degree, clustering coefficient, node coloring, etc.) and then use a kernel function to measure the similarity between graphs based on these features. On the other hand, deep learning-based methods leverage neural networks to learn representations of graphs, allowing for more flexible and powerful classification models, which may not involve hand-crafted features or some possibly biased assumptions about the graph structure, as it learns the features directly from the data. In this context, Graph Representation Learning (GRL), also known as Geometric Deep Learning [5], has emerged as a powerful framework for learning representations of graph-structured data. The basic idea is that — as common in most deep learning approaches — the model learns a mapping function from the input space (i.e., the graph) to a high-dimensional embedding space (i.e., a latent space where a graph $G_{i}$ is represented as a vector $h_{G_{i}}$) using a neural network. Such representation should preserve some useful properties of the original input data, and should be useful for the downstream task like, but not limited to, graph classification.
@@
The most popular approach to learn representations of graphs are Graph Neural Networks (GNNs) [19, 4, 39], based on a message-passing mechanism that allows each node to exchange information with its neighbors and to build its own representation based on a trainable function of the information received from its neighbors. That is, GNNs iteratively update the representation of each node by aggregating incoming messages from its neighbors, which is, in turn, sent to its own neighbors in the next iteration to update their representations. A plethora of GNNs have been proposed in the literature, each with its own architecture and design choices. For instance, Graph Convolutional Networks (GCNs) [27] use a convolutional operator (i.e., a shared filter) to aggregate information from neighboring nodes, GraphSAGE [20] extends this idea by using a sampling strategy and sampling a fixed-size neighborhood to allow the model to scale to large graphs, while Graph Attention Networks (GATs) [veličković2018graphattentionnetworks, 14] use an attention mechanism to weigh the importance of neighboring nodes when aggregating their features, allowing the model to focus on the most relevant ones for the task, and Graph Isomorphism Networks (GIN) [41] are inspired by the Weisfeiler-Lehman graph isomorphism test and use a neural network (e.g., a multi-layer perceptron) to learn a representation of the graph based on its structure.
@@
From node representations, graph-level representations (i.e., the graph embedding) can be obtained by pooling the representation of its most elementary components (i.e., nodes) using a permutation-invariant pooling function like global pooling functions (e.g., mean, max, sum) or by leveraging more sophisticated techniques such as attention mechanisms or hierarchical aggregation [42, 38]. Specifically, a Geometric Deep Learning (GDL) model is usually composed of a series of GNN layers, interspersed with non-linear activation functions, followed by the pooling layer. That is, $h_{G}=\operatorname{pool}(h_{v_{1}},h_{v_{2}},\ldots,h_{v_{n}})$, where $h_{G}$ is the graph-level representation, $h_{v_{i}}$ is the representation of node $v_{i}$, and $\operatorname{pool}$ is a permutation-invariant pooling function.
@@
Once the graph-level representation is obtained, it can be used for various tasks (including, of course, graph classification) using classic ML techniques, such as fully connected layers, support vector machines, or other classifiers. In this sense, similarity between two graphs can be defined as the distance between their respective representations in the embedding space, which is often learned by the GDL model for a specific task, such as graph classification. More formally, given two graphs $G_{i}$ and $G_{j}$, they are similar if their representations $h_{G_{i}}$ and $h_{G_{j}}$ are close in the embedding space. In this context, graph similarity computation [24, 32] has seen significant advancements in recent years and has proven to be highly beneficial across various domains, but has not yet been tested in the field of graph generation extensively. A method for improving graph generation evaluation using a classifier was proposed in the literature by Liu et al. [31], although it was demonstrated on a limited dataset and with restricted use of Graph Generative Models. Specifically, they trained both the Graph Generative Models and a Graph Isomorphism Network (GIN) classifier on small graph datasets such as Community-small, Grid-small, and Ego-small.While this work represents a step forward, it has several limitations, including the use of small datasets and the limited evaluation of the generated graphs. Here, we address these limitations by proposing a more comprehensive evaluation methodology that leverages larger datasets that include both real-world and synthetic graphs, by employing a more sophisticated model architecture and training methodology, and by providing a more thorough evaluation of the generated graphs.